OpenStudy (anonymous):

How would I do the line integral for Fdr where F=2xyi+x^2j and the line is described by r(t)=(t^2-2t)i+(tsin(t-2))j over 0

OpenStudy (anonymous):

\[\int\limits Fdr, F=<2xy,x^2>, C: r(t)+<t^2-2t, tsin(t-2)>, 0<t<2\]

OpenStudy (anonymous):

actually, is this green's theorem, since F is only 2 dimensions?

OpenStudy (anonymous):

Would it come out to zero?

OpenStudy (anonymous):

I guess in general I'm having trouble understanding how to construct the surface integral for Stokes. I know how to find the curl, but I'm not sure what I'm supposed to be dotting it with (dS? a normal vector? what is dS?)

OpenStudy (amistre64):

F(r(t)).r'(t) sounds familiar

OpenStudy (amistre64):

im not remembering the specific for stokes to be sure what to do on that end of it

OpenStudy (anonymous):

Thanks for the reply. I believe F(r(t).r'(t)) is if i want to integrate it directly.

OpenStudy (amistre64):

F= (2xy,x^2) x=t^2-2t ; x' = 2t - 2 y=t sin(t-2) ; y' = t cos(t-2) + sin(t-2) F(r(t))= 2 t sin(t-2) (t^2-2t) ; (t^2-2t)^2 2t - 2 ; t cos(t-2) + sin(t-2) --------------------------------------- whatever the dot is .............................

OpenStudy (amistre64):

correct, if you just want to turn it all into one big sum of a quazihoots yes

OpenStudy (amistre64):

if you want a readable explanation and example of Greens, this site is always good for me http://tutorial.math.lamar.edu/Classes/CalcIII/GreensTheorem.aspx

OpenStudy (anonymous):

:) yeah, it gets really ugly. I'm pretty sure the problem is engineered for Stokes, just haven't gotten a good grasp on how to apply it.

OpenStudy (anonymous):

hey, i'll check it out. thank you.

OpenStudy (amistre64):

that should have stokes on the site somewhere as well; if not theres a patrickJMT or another on youtube that is also pretty easy to follow for these things

OpenStudy (amistre64):

its just been to long past and too few practices to recollect these for me :)

OpenStudy (anonymous):

hey, that's some good stuff. I appreciate the help.

OpenStudy (amistre64):

good luck :)